Assume there are two planes Π1 and Π2 for which there exists a line of intersection L.
a) Suppose you find two different parametric forms for L using two different methods. How can you prove they are the same line?
b) Let there be vectors n1 and n2 that are normals to Π1 and Π2 respectively. If m = n1×n2 (cross product) and m is parallel to L, give a geometric description of this result.
1
Expert's answer
2019-07-03T14:31:36-0400

a) If two different equations of the line are


"(x_1, y_1, z_1)+t_1(a, b, c)\ \ \ (1)"

and


"(x_2, y_2, z_2)+t_2(d, e, f)\ \ \ (2)"

where t1 and t2 are parameters.

Since there is only one line containing given point and parallel to given non-zero vector, it is enough to prove that both lines contains the same point and their directional vectors are parallel and therefore have proportional coordinates:


"a\/d=b\/e=c\/f"


If equations (1) and (2) represent the same line, point (x2, y2, z2) should satisfy equation (1)


"(x_1, y_1, z_1)+t_1(a, b, c)=(x_2, y_2, z_2)"

so system


"x_1+at_1=x_2,\ny_1+bt_1=y_2,\nz_1+ct_1=z_2"

should have unique solution t1.


In general to prove that two equations represent the same curve it is enough to find a bijection between t1 and t2, so that corresponding point of curves are the same.

In case of ф straight line if a bijection exists, then it can be found from the equation


"x_1+at_1=x_2+dt_2"

or

"t_2=(x_1-x_2)\/d+at_1\/d"

and should satisfy equations


"y_1+bt_1=y_2+et_2\ and\ z_1+ct_1=z_2+ft_2"

b)The line of intersection L belongs to two planes Π1 and Π2, hence L is orthogonal to normal vectors n1 and n2.

Since planes Π1 and Π2 are not parallel, their normal vectors are not parallel either.

A direction of vector orthogonal to two given linearly independent vectors is uniquely determined and the vector m = n1×n2  is orthogonal to vectors n1 and n2, hence a directional vector of L is parallel to m.

Another geometric description is as follows:

Vector m = n1×n2 is orthogonal to vectors n1 and n2, hence m is parallel to Π1 and Π2, therefore m is parallel to the line L of intersection of planes.